OpenStudy (anonymous):

I need help understanding somethin pls help Ok, so when you have a fraction like this -(-3/-5), does it turn into -(3/5)?

OpenStudy (anonymous):

Because someone told me that when there are negative signs in a parenthesis it turns positive

OpenStudy (calculusxy):

do you mean: \(\huge-(\frac{-3}{-5})\)

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

since the numerator is a negative and the denominator is also negative remember the rule: \(\huge - \div - = +\) so \(-3 \div -5\) is no exception. so this will be \(\large -(\frac{3}{5})\). Since the negative applies to the whole fraction (since it's in parenthesis), it will be \(\large -\frac{3}{5}\) .

OpenStudy (anonymous):

Ok thanks for explaining, I have another question, how does -(-3/5) turn into only 3/5? Isn't it the same rule? Like when there any negatives in the parenthesis it turns into a positive?

OpenStudy (calculusxy):

well: \(\huge - \times - = +\) it basically like you have \(\large -1 \times -\frac{3}{5} = \frac{3}{5}\)

OpenStudy (anonymous):

uh still don't get it

OpenStudy (skullpatrol):

think of -a meaning the "opposite" of a then -(-a) = a

OpenStudy (skullpatrol):

-2 is the opposite of 2 so -(-2) = 2 the opposite of the opposite of 2 is 2

OpenStudy (calculusxy):

usually in front of parenthesis you would have a 1 (well that's how i think of it). if we think of this in this problem we would have: \(\huge \color{red}{-1}(-\frac{3}{5})\) how would you evaluate this?

OpenStudy (anonymous):

um

OpenStudy (anonymous):

-1(3/5)?

OpenStudy (anonymous):

hello?

OpenStudy (skullpatrol):

yes :-)

OpenStudy (skullpatrol):

$$\Huge -a=-1a$$

OpenStudy (anonymous):

so uh it would be -(3/5)?

OpenStudy (skullpatrol):

what is the opposite of 3/5?

OpenStudy (anonymous):

-3/5

OpenStudy (skullpatrol):

that is the simplified answer :D

OpenStudy (anonymous):

uh thank you

OpenStudy (skullpatrol):

Thanks for trying to learn :-)

OpenStudy (anonymous):

you can leave now I don't need any help right now ^_^